OpenStudy (anonymous):

Amy buys a pair of running shoes for 57.37durring a 40% off sale what was the original price?

jimthompson5910 (jim_thompson5910):

Let x = original price The word problem then translates to x - 0.4x = 57.37 Now solve for x x - 0.4x = 57.37 0.6x = 57.37 x = 57.37/0.6 x = 95.6167 Now round to the nearest cent to get x = 95.62 So the original price was $95.62

OpenStudy (anonymous):

I know you used more than a program to figure that out. ha!

jimthompson5910 (jim_thompson5910):

lol thx

OpenStudy (anonymous):

and besides accessing you is like accessing a program so what's the beef?

jimthompson5910 (jim_thompson5910):

who knows lol